OpenStudy (anonymous):

25. Find all the zeroes of the equation. 27x2 – 324 = –x4 (1 point)–3, –6i 3, 6i 3, –3, 6i, –6i 3, –3, 6i, 0

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so rewrite the equation \[x^4 + 27x^2 - 324 = 0\]

OpenStudy (campbell_st):

ok... so this is an equation that can be treated as a quadratic the factors of -324 that add to 27 are 26 and - 9 so you can factorise it to \[(x^2 - 9)(x^2 + 36) = 0\] so you need to solve \[x^2 - 9 = 0\] and \[x^2 + 36 = 0\] one zero is rational and one is complex hope this helps

OpenStudy (anonymous):

so my answer is C right?

OpenStudy (campbell_st):

yep... looks like it

OpenStudy (anonymous):

thank u so much!!!